OpenStudy (anonymous):

.Camille wants to buy a new coat and gloves for her winter ski trip. She finds the perfect coat for $79.50 and matching gloves for $14.50. When she gets to the cash register she finds out that all coats are on sale for 20% off. The sales tax rate in Camille's city is 6%. What is the total amount Camille has to pay for the two items, including sales tax?

OpenStudy (zeig_101):

1)subtract 20%of 79.50 2)79.50(minus the 20%)+14.50 3)now add the 6% of total

OpenStudy (anonymous):

THanks ! :D

OpenStudy (zeig_101):

Post your work here so i can check it;)

OpenStudy (anonymous):

ok well what is the money u have left

OpenStudy (anonymous):

1)subtract 20% of 79.50= -79.3 2)79.50(minus the 20%)+14.50= -64.8 3)now add the 6% of total = 3.89 My Work :)

OpenStudy (zeig_101):

Now complete the problem with these numbers

OpenStudy (anonymous):

= -68.69?

OpenStudy (zeig_101):

no, sorry, rework your problem

OpenStudy (anonymous):

oh ok

OpenStudy (zeig_101):

you take 20% of 79.50, then subtract is from 79.50

OpenStudy (anonymous):

Oh..woops XD

OpenStudy (zeig_101):

I had a hard time getting used to this kind of problem as well, so don't worry about it.

OpenStudy (anonymous):

good..i thought it was only me lol

OpenStudy (anonymous):

is it -45.21? thats what i got now

OpenStudy (zeig_101):

here let me put up a bit of the work

OpenStudy (zeig_101):

79.50-20% of 79.50 new number+the 14.50 for the gloves then take 6% of the total and add it onto the total Does this help any?

OpenStudy (anonymous):

yea..im going to try again like that

OpenStudy (anonymous):

i got 82.79

OpenStudy (zeig_101):

79.50-15.90=63.60 63.60+14.50=78.10 78.10+4.686 4.686 rounded is 4.69 78.10+4.69=82.79

OpenStudy (zeig_101):

you are correct!

OpenStudy (anonymous):

:D thank you! i gave you a medal, thanks so much though!'

OpenStudy (zeig_101):

And I will give you a medal for being a good sport!

OpenStudy (anonymous):

:)

OpenStudy (zeig_101):

You may now close the question.